OpenStudy (anonymous):

For a right triangle that has legs a and b and hypotenuse c, find the length of the side not given here. a= 18 inches c=30 inches

OpenStudy (gir_lover_♥):

Use the pythagrium thereom

OpenStudy (gir_lover_♥):

a^2+b^2=c^2

OpenStudy (gir_lover_♥):

18^2+ ?=c^2

OpenStudy (gir_lover_♥):

324+ ? = 900 Subtract 900-324

OpenStudy (gir_lover_♥):

Then u divid that number to simplify it